# Extrapolation and Factorization of matrix weights

### v3

@@ Abstract.
The first author was partially supported by NSF grant DMS-1956395. The second author was supported by research funds from the Dean of the College of Arts & Sciences, the University of Alabama, and is currently partially supported by a Simons Foundation Travel Support for Mathematicians Grant. The authors would like to thank Fedor Nazarov for an inspiring lecture on matrix weights at the University of Missouri over two decades ago, which led to the formulation of Proposition 8.8.
@@ 1. Introduction
Given this background, we can now describe our main results. To do so we must first introduce a change in notation. For a number of reasons connected to our proofs, we have chosen to write matrix weighted norms in the form
$\bigg{(}\int_{\mathbb{R}^{n}}|W(x){f}(x)|^{p}\,dx\bigg{)}^{\frac{1}{p}}.$
@@
Let $T$ be a sublinear operator. Suppose that for some $p_{0}$, $1\leq p_{0}\leq\infty$, there exists an increasing function $K_{p_{0}}$ such that for every $W_{0}\in\mathcal{A}_{p_{0}}$,
@@
Then for all $p$, $1<p<\infty$, and for all $W\in\mathcal{A}_{p}$,
@@ 10. An application of Rubio de Francia extrapolation
Let $T$ be a sublinear scalar-valued operator. We assume that the extension of $T$ to vector-valued functions, which is given for ${f}=(f_{1},\ldots,f_{d})^{t}$ by applying it to each coordinate $T{f}=(Tf_{1},\ldots,Tf_{d})^{t}$, fulfills the hypothesis of Theorem 1.4. That is, for some $p_{0}$, $1\leq p_{0}\leq\infty$, there exists an increasing function $K_{p_{0}}$ such that for every $W_{0}\in\mathcal{A}_{p_{0}}$ we have (1.5). By the scalar-valued extrapolation theorem, this implies that $T$ is bounded on the scalar weighted spaces $L^{p}(\mathbb{R}^{n},w)$, for any $w\in\mathcal{A}_{p}$, $1<p<\infty$. To apply Theorem 9.1, we need to construct a family $\mathcal{F}$ of pairs $(Tf,f)$, such that given any $p$, $1<p<\infty$, and $W\in\mathcal{A}_{p}$, then for any pair $(Tf,f)\in\mathcal{F}$, we have $\|Tf\|_{L^{p}(\mathbb{R}^{n},W)}<\infty$.
@@
then for each $1<p<\infty$ we can apply the conclusion of Theorem 9.1 to the family $\mathcal{F}$. This gives us the desired inequality (1.6) for all $f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$. By [18, Propositions 3.6, 3.7] we have that $L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$ (as well as $C_{c}^{\infty}(\mathbb{R}^{n},\mathbb{R}^{d})$) are dense in $L^{p}(\mathbb{R}^{n},W)$ for any matrix weight $W$ and $1\leq p<\infty$. Therefore, by a standard approximation argument we get the inequality (1.6) for all $f\in L^{p}(\mathbb{R}^{n},W)$. ∎
Take $\Omega\in L^{\infty}(S^{n-1})$ with $\Omega||_{\infty}\leq 1$ and vanishing integral on the unit sphere $S^{n-1}$ in $\mathbb{R}^{n}$. For each $0<\delta<1$, define truncated rough singular integral by
@@
$T_{\Omega}^{\natural}f(x)=\sup_{\delta>0}|T_{\Omega,\delta}f(x)|.$
@@
(10.1) $\|\sup_{\delta>0}|WT_{\Omega,\delta}f|\|_{L^{2}(\mathbb{R}^{n},W)}\leq C[W]_{\mathcal{A}_{2}}^{5}\|f\|_{L^{2}(\mathbb{R}^{n},W)}.$
@@
(10.2) $\|\sup_{\delta>0}|WT_{\Omega,\delta}f|\|_{L^{p}(\mathbb{R}^{n},W)}\leq C[W]_{\mathcal{A}_{p}}^{5\max\{\frac{p}{2},\frac{p^{\prime}}{2}\}}\|f\|_{L^{2}(\mathbb{R}^{n},W)}.$
@@
Following the proof of Theorem 1.4, we construct an appropriate family $\mathcal{F}$ of extrapolation pairs. Fix $f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$; then $f\in L^{p}(W)$ for all $p$ and $W\in\mathcal{A}_{p}$. Since $T_{\Omega}^{\natural}$ is bounded on $L^{p}(\mathbb{R}^{n},\mathbb{R}^{d})$, we have that for almost every $x\in\mathbb{R}^{n}$
$\sup_{\delta>0}|T_{\Omega,\delta}f(x)|<\infty.$
@@
$\sup_{\delta>0}|W(x)T_{\Omega,\delta}f(x)|<\infty$
@@
$\sup_{\delta>0}|W(x)T_{\Omega,\delta}f(x)|\leq 2|W(x)T_{\Omega,\delta_{x}}f(x)|.$
Let $g(x)=T_{\Omega,\delta_{x}}f(x)$. Since $T_{\Omega}^{\natural}$ is bounded on $L^{p}(\mathbb{R}^{n},w)$ for all scalar $w\in A_{p}$, we have that
@@
$\int_{\mathbb{R}^{n}}|V(x)g(x)|^{2}\,dx\leq\int_{\mathbb{R}^{n}}\big{(}\sup_{\delta>0}|V(x)T_{\Omega,\delta}f(x)|\big{)}^{2}\,dx\leq C[V]_{\mathcal{A}_{2}}^{5}\|f\|_{L^{2}(\mathbb{R}^{n},V)}.$
Therefore, if we let $\mathcal{F}=\{(g,f):f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})\}$, where $g(x)=T_{\Omega,\delta_{x}}f(x)$, we have that this family satisfies the hypotheses of Theorem 9.1. Hence, we can conclude that (10.2) holds for all $f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$. Since the operator $f\mapsto\sup_{\delta>0}|WT_{\Omega,\delta}f|$ is sublinear, by an approximation argument the inequality holds for all $f\in L^{p}(\mathbb{R}^{n},W)$. ∎

### v4

@@ Abstract.
The first author was partially supported by NSF grants DMS-1956395 and DMS-2349756. The second author was supported by research funds from the Dean of the College of Arts & Sciences, the University of Alabama, and is currently partially supported by a Simons Foundation Travel Support for Mathematicians Grant. The authors would like to thank Fedor Nazarov for an inspiring lecture on matrix weights at the University of Missouri over two decades ago, which led to the formulation of Proposition 8.8.
@@ 1. Introduction
Given this background, we can now describe our main results. To do so we must first introduce a change in notation. For a number of reasons connected to our proofs, we have chosen to write a matrix weighted norm of a function $f:\mathbb{R}^{n}\to\mathbb{R}^{d}$ in the form
$f||_{L^{p}(\mathbb{R}^{n},W)}=\bigg(\int_{\mathbb{R}^{n}}|W(x){f}(x)|^{p}\,dx\bigg)^{\frac{1}{p}}.$
@@
Given an operator $T$, suppose that for some $p_{0}$, $1\leq p_{0}\leq\infty$, there exists an increasing function $K_{p_{0}}$ such that for every $W_{0}\in\mathcal{A}_{p_{0}}$,
@@
Then for all $p$, $1<p<\infty$, for all $W\in\mathcal{A}_{p}$, and for all $f\in L^{\infty}_{c}(\mathbb{R}^{n})$,
@@
Moreover, if $T$ is linear, it has a continuous extension to all $f\in L^{p}(\mathbb{R}^{n},W)$ that satisfies the same bound.
@@
Remark 1.8.
By [18, Propositions 3.6, 3.7] we have that $L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$ is dense in $L^{p}(\mathbb{R}^{n},W)$ for any matrix weight $W$ and $1\leq p<\infty$. In Theorem 1.4, the set $L^{\infty}_{c}(\mathbb{R}^{n})$ can be replaced by any collection which is dense in $L^{p}(\mathbb{R}^{n},W)$ and contained in every scalar weighted space $L^{p}(\mathbb{R}^{n},w)$. If $T$ is not linear, the problem of proving the continuous extension exists is more delicate, as is the problem of showing this abstract extension agrees with the original operator. We consider a specific example in Theorem 10.1 below.
@@ 10. An application of Rubio de Francia extrapolation
Let $T$ be a scalar-valued operator. We assume that the extension of $T$ to vector-valued functions, which is given for ${f}=(f_{1},\ldots,f_{d})^{t}$ by applying it to each coordinate $T{f}=(Tf_{1},\ldots,Tf_{d})^{t}$, fulfills the hypothesis of Theorem 1.4. That is, for some $p_{0}$, $1\leq p_{0}\leq\infty$, there exists an increasing function $K_{p_{0}}$ such that for every $W_{0}\in\mathcal{A}_{p_{0}}$ we have (1.5). If we take any scalar weight $w_{0}\in\mathcal{A}_{p_{0}}$ we can define $W_{0}$ to be the diagonal matrix with copies of $w_{0}$ on the diagonal, so $T$ is bounded on $L^{p_{0}}(\mathbb{R}^{n},w_{0})$. Hence, the scalar-valued extrapolation theorem implies that $T$ is bounded on the scalar weighted spaces $L^{p}(\mathbb{R}^{n},w)$, for any $w\in\mathcal{A}_{p}$, $1<p<\infty$. To apply Theorem 9.1, we need to construct a family $\mathcal{F}$ of pairs $(Tf,f)$, such that given any $p$, $1<p<\infty$, and $W\in\mathcal{A}_{p}$, then for any pair $(Tf,f)\in\mathcal{F}$, we have $\|Tf\|_{L^{p}(\mathbb{R}^{n},W)}<\infty$.
@@
then for each $1<p<\infty$ we can apply the conclusion of Theorem 9.1 to the family $\mathcal{F}$. This gives us the desired inequality (1.6) for all $f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$.
Now suppose that $T$ is linear. Take any $f\in L^{p}(\mathbb{R}^{n},W)$. Let $\{f_{j}\}_{j=1}^{\infty}$ be a sequence in $L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$, which converges to $f$ in $L^{p}(\mathbb{R}^{n},W)$ norm. Since $T$ is linear, we have $Tf_{j}-Tf_{k}=T(f_{j}-f_{k})$, and by (1.6), $\{Tf_{j}\}_{j=1}^{\infty}$ is a Cauchy sequence in $L^{p}(\mathbb{R}^{n},W)$. Since the space $L^{p}(\mathbb{R}^{n},W)$ is complete, this sequence converges. Define $Tf$ to be the limit. Thus, we get the inequality (1.6) for all $f\in L^{p}(\mathbb{R}^{n},W)$. ∎
Take $\Omega\in L^{\infty}(S^{n-1})$ with $\Omega||_{\infty}\leq 1$ and vanishing integral on the unit sphere $S^{n-1}$ in $\mathbb{R}^{n}$. For each $0<\delta<1$, define the truncated rough singular integral by
@@
$T_{\Omega}^{\natural}f(x)=\sup_{0<\delta<1}|T_{\Omega,\delta}f(x)|.$
@@
(10.1) $\|\sup_{0<\delta<1}|WT_{\Omega,\delta}f|\|_{L^{2}(\mathbb{R}^{n})}\leq C[W]_{\mathcal{A}_{2}}^{5}\|f\|_{L^{2}(\mathbb{R}^{n},W)}.$
@@
(10.2) $\|\sup_{0<\delta<1}|WT_{\Omega,\delta}f|\|_{L^{p}(\mathbb{R}^{n})}\leq C[W]_{\mathcal{A}_{p}}^{5\max\{\frac{p}{2},\frac{p^{\prime}}{2}\}}\|f\|_{L^{p}(\mathbb{R}^{n},W)}.$
@@
Fix $p$, $1<p<\infty$. Following the proof of Theorem 1.4, we construct an appropriate family $\mathcal{F}$ of extrapolation pairs. Fix $f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$; then $f\in L^{p}(W)$ for all $p$ and $W\in\mathcal{A}_{p}$. Since $T_{\Omega}^{\natural}$ is bounded on $L^{p}(\mathbb{R}^{n},\mathbb{R}^{d})$, we have that for almost every $x\in\mathbb{R}^{n}$
$\sup_{0<\delta<1}|T_{\Omega,\delta}f(x)|<\infty.$
@@
$\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f(x)|<\infty$
@@
$\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f(x)|\leq 2|W(x)T_{\Omega,\delta_{x}}f(x)|.$
Let $g(x)=T_{\Omega,\delta_{x}}f(x)$; it is straightforward to show that we can choose $\delta_{x}$ measurably, so $g$ is a measurable function. Since $T_{\Omega}^{\natural}$ is bounded on $L^{p}(\mathbb{R}^{n},w)$ for all scalar $w\in\mathcal{A}_{p}$ [28], by Corollary 6.7 we have that
@@
$\bigg(\int_{\mathbb{R}^{n}}|V(x)g(x)|^{2}\,dx\bigg)^{\frac{1}{2}}\leq\bigg(\int_{\mathbb{R}^{n}}\big(\sup_{0<\delta<1}|V(x)T_{\Omega,\delta}f(x)|\big)^{2}\,dx\bigg)^{\frac{1}{2}}\leq C[V]_{\mathcal{A}_{2}}^{5}\|f\|_{L^{2}(\mathbb{R}^{n},V)}.$
Therefore, if we let $\mathcal{F}=\{(g,f):f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d}),g(x)=T_{\Omega,\delta_{x}}f(x)\}$, we have that this family satisfies the hypotheses of Theorem 9.1. Hence, it follows that (10.2) holds for all $f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$.
Now fix $p$, $1<p<\infty$, and $W\in\mathcal{A}_{p}$. Since the operator $T_{\Omega,\delta}$ is linear, the operator $T$, which is defined initially for $f\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$ by
$Tf(x)=\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f(x)|$
is sublinear. Arguing as we did above in the proof of Theorem 1.4, we can show that $T$ has a continuous extension to all $f\in L^{p}(\mathbb{R}^{n},W)$. Indeed, take any $f\in L^{p}(\mathbb{R}^{n},W)$. Let $\{f_{j}\}_{j=1}^{\infty}$ be a sequence in $L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$, which converges to $f$ in $L^{p}(\mathbb{R}^{n},W)$ norm. Since $T$ is sublinear, we have $|Tf_{j}-Tf_{k}|\leq|T(f_{j}-f_{k})|$, and by (10.2), $\{Tf_{j}\}_{j=1}^{\infty}$ is a Cauchy sequence in $L^{p}(\mathbb{R}^{n})$, so this sequence converges. Define $Tf$ to be the limit. Clearly, $T:L^{p}(\mathbb{R}^{n},W)\to L^{p}(\mathbb{R}^{n})$ is bounded.
We will now show that for all $f\in L^{p}(\mathbb{R}^{n},W)$,
(10.3) $\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f(x)|\leq Tf(x)$
almost everywhere. Fix such an $f$ and for each $k\in\mathbb{N}$, define $f_{k}\in L^{\infty}_{c}(\mathbb{R}^{n},\mathbb{R}^{d})$ by
$f_{k}(x)=f(x)\min\bigg(1,\frac{k}{|f(x)|}\bigg)\chi_{B(0,k)}(x).$
Then we have that $|f_{k}(x)|\leq|f(x)|$ and $f_{k}\to f$ pointwise a.e. Moreover, since $f$ and $f_{k}$ are parallel vectors, $|W(x)f_{k}(x)|\leq|W(x)f(x)|$. Hence, by the dominated convergence theorem, $f_{k}\to f$ in $L^{p}(\mathbb{R}^{n},W)$. Therefore, we have that $\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f_{k}|$ converges to $Tf$ in $L^{p}(\mathbb{R}^{n})$. By passing to a subsequence, we then have that for almost every $x\in\mathbb{R}^{n}$,
(10.4) $Tf(x)=\lim_{k\to\infty}\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f_{k}(x)|.$
Now fix $\delta$, $0<\delta<1,$ and let $B=B(0,N)$, where $N>\delta^{-1}$. For brevity, in the definition of $T_{\Omega,\delta}$ we write $x-y]^{\prime}=(x-y)|x-y|^{-1}$. Then we have that
$\int_{B}|W(x)T_{\Omega,\delta}f(x)|^{p}\,dx$ $=\int_{B}\bigg|W(x)\int_{\delta<|x-y|<\delta^{-1}}\frac{\Omega([x-y]^{\prime})}{|x-y|^{n}}f(y)\,dy\bigg|^{p}\,dx$
$\leq\delta^{-np}\int_{B}\bigg(\int_{|x-y|<\delta^{-1}}|W(x)W^{-1}(y)W(y)f(y)|\,dy\bigg)^{p}\,dx$
$\leq\delta^{-np}\int_{2B}\bigg(\int_{2B}|W(x)W^{-1}(y)|^{p^{\prime}}_{\operatorname{op}}\,dy\bigg)^{\frac{p}{p^{\prime}}}\,dx\bigg(\int_{\mathbb{R}^{n}}|W(y)f(y)|^{p}\,dy\bigg)$
$\leq C[W]_{\mathcal{A}_{p}}^{p}|B|^{p}\|f\|^{p}_{L^{p}(\mathbb{R}^{n},W)};$
the last inequality holds by applying the $\mathcal{A}_{p}$ condition with balls instead of cubes. Since $N$ can be made arbitrarily large, we get that for all $\delta$, $WT_{\Omega,\delta}f\in L^{p}_{loc}(\mathbb{R}^{n})$, and so $|WT_{\Omega,\delta}f(x)|<\infty$ a.e.
Moreover, again since $f$ and $f_{k}$ are parallel vectors, for a.e. $x,\,y\in\mathbb{R}^{n}$
$|W(x)\Omega([x-y]^{\prime})|x-y|^{-n}f_{k}(y)|\leq|W(x)\Omega([x-y]^{\prime})|x-y|^{-n}f(y)|.$
Therefore, by the dominated convergence theorem, we have that for every $\delta$ and for almost every $x$,
(10.5) $W(x)T_{\Omega,\delta}f(x)=\lim_{k\to\infty}W(x)T_{\Omega,\delta}f_{k}(x).$
Fix an $x$ such that both (10.4) and (10.5) hold. Consequently,
$\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f(x)|\big|$ $=\sup_{0<\delta<1}\lim_{k\to\infty}|W(x)T_{\Omega,\delta}f_{k}(x)|$
$\leq\lim_{k\to\infty}\sup_{0<\delta<1}|W(x)T_{\Omega,\delta}f_{k}(x)|=Tf(x).$
This proves (10.3) and the boundedness of $T:L^{p}(\mathbb{R}^{n},W)\to L^{p}(\mathbb{R}^{n})$ yields (10.2). ∎